OpenStudy (anonymous):

An athlete can burn 10 calories per minute by running and 4 calories per minute by walking. If the athlete burns 450 calories in 60 minutes, how long is spent on each activity?

OpenStudy (anonymous):

5a+2b=225, a+b=60 implies a=35 and b=25 minutes. Running 35 minutes and walking 25 minutes

OpenStudy (anonymous):

How did you get the 5a+2b+225? a+b=60 I get that. although I was figuring a+b=450.

OpenStudy (anonymous):

It looks like the 10 seconds and the 4 seconds and the total calories was divided into half.